# An interpretable generative multimodal neuroimaging-genomics framework for decoding Alzheimer’s disease

### v1

@@ 2.2 Preprocessing and feature engineering
The sMRI volumes preprocessing included tissue segmentation in Gray Matter (GM), White matter and Cerebrospinal fluid (CSF) using the modulated normalization algorithm. Only the GM volume was considered for this study, to which a smoothing using a Gaussian kernel (FWHM = $6$mm) was applied. Quality control (QC) included discarding images that exhibited low correlation with individual and/or group level masks. The full preprocessed GM volume was used as input for the sMRI channel resulting in an input size of $121\times 145\times 121$ for each subject.
@@
The rs-fMRI data was preprocessed using the statistical parametric mapping toolbox (SPM12, http://www.fil.ion.ucl.ac.uk/spm/) including rigid body motion correction to correct subject head motion, slice-timing correction, warping to the standard MNI space using the EPI template, resampling to ${(3mm)}^{3}$ isotropic voxels, and smoothing using a Gaussian kernel (FWHM = 6mm), following the preprocessing proposed in [47]. QC procedure was the same as for sMRI. Fifty-three independent components (ICs) covering the whole brain were extracted using spatially constrained ICA with the Neuromark_fMRI_1.01.01.01.0 template (available in the GIFT software; http://trendscenter.org/software/gift). For each subject, a correlation matrix was then created computing the Pearson correlation between the ICs time courses, resulting in a $53$x53535353 static functional network connectivity (sFNC) matrix. This was divided in $7$ RSNs, named: (i) Sub-cortical network (SC); (ii) Auditory network; (iii) Sensorimotor network (SM); (iv) Visual network (VI); (v) Cognitive-control network (CC); (vi) Default-mode network (DM); and, (vii) Cerebellar network (CB) (please, refer to Supplementary Table 2 for more information). The upper triangular matrix was then vectorized resulting in an input vector of 1378 features for each subject.
@@
Moving to the genomics data, pre-imputation QC was performed to remove Single Nucleotide Polymorphisms (SNPs) with minor allele frequency (MAF) $<0.05$, call rate $<0.98$, and Hardy Weinberg Equilibrium $<10^{-3}$ (details in https://www.synapse.org/#!Synapse:syn2290704/wiki/64710). Imputation was performed with the same reference panel. Only the SNPs with imputation $r^{2}>0.4$ were included. Linkage disequilibrium (LD)-based pruning with $r^{2}=0.8$ in a window of $50$kb was applied, yielding $445838$ SNPs for further analyses [48]. A feature selection leveraging on the genome-wide association study (GWAS) on AD conducted by [17] was carried on to include all the relevant SNPs having GWAS p-values less than $1$e-04. A total of $565$ polymorphisms was selected and a value between 0 and 2 was assigned based on the presence of mutated alleles. In detail defining wild and mutated alleles respectively as W and M a score of 00 means wild homozygous alleles (W/W), a score of $1$ indicates heterozygous alleles (W/M) and a score of $2$ defines mutated homozygous alleles (M/M).
2.3 Classification tasks
Following the aim of this study, the proposed architecture was devised with a twofold classification aim: (1) AD detection, that is the differentiation of AD and CN subjects, also referred as Task 1 and (2) MCI conversion prediction, that is the stratification of MCIc and MCInc patients, also referred as Task 2. In detail, the network was trained and tested on Task 1 and subsequently used to solve Task 2 allowing to assess its ability in discriminating different stages of disease and also in capturing and highlighting shared patterns across the different categories.
2.4 Framework architectures
The proposed framework builds on our previous work [30] and is shown in Figure 1. In detail, the architecture for disease detection consists of three modules: i) Feature reduction module, which performs a CNN-based feature extraction to derive a lower dimensionality latent space, separately for each input channel ii) Generative module that, in the eventuality of missing modalities, actuates a generative process in the latent space transferring the knowledge from one domain to another; iii) Data fusion & classification module that fuses the latent features obtained for the three modalities and then performs the classification. Post-hoc interpretability analysis was then carried on in order to retrieve feature contributions to the classification task. The three modules will be detailed in the following paragraphs.
@@ 2.4 Framework architectures
Figure 1: Schematic representation of the proposed framework. Given three input modalities, structural Magnetic Resonance Imaging (sMRI), functional MRI (fMRI) and Single Nucleotide Polymorphisms (SNPs), the framework is articulated in three different modules: i) A feature reduction module where the input channels are transformed in their low dimensionality latent representations; ii) A generative module where, if necessary, the missing modalities are imputed transferring the knowledge from one domain to another; iii) A data fusion & classification module where the latent features are fused together and then classified using a Multilayer Perceptron. Finally, post-hoc interpretability analysis is performed relying on Integrated Gradients (IG) method for feature attribution derivation.
@@
The feature reduction module consists of three different CNNs, one for each input channel, leading to a latent low dimensionality representation consisting of $100$ latent features for each channel. The sMRI channel was analyzed through a 3D CNN defined by three successive convolutional blocks, each composed of two consecutive convolutional layers (filter sizes of $3\times 3\times 3$ for the first four convolutional layers and $2\times 2\times 2$ for the last two, number of filters: $64$, $64$, $64$, $128$, $128$, and $128$, respectively, for the six convolutional layers) and one max-pooling layer, followed by four fully-connected layers (FCLs) (number of nodes: $1536$, $768$, $384$, and $192$, respectively).
The rs-fMRI channel was analyzed through a $1$D CNN, composed by four convolutional layers (filter sizes $5\times 5\times 5$, number of filters: $64$, $128$, $128$, and $128$, respectively), two max-pooling layers and two FCLs (number of nodes: $384$ and $100$, respectively).
Finally, for genetic data, a $1$D CNN was employed, consisting of three convolutional blocks, each including a sequence of convolutional layers (filter sizes $3\times 3\times 3$, number of filters: $64$, $64$, and $128$, respectively, for the three convolutional layers), batch-normalization and max-pooling, followed by three FCLs (number of nodes: $1024$, $512$, and $128$, respectively).
The ReLU activation function [49] was used for all the layers of these three CNNs.
@@
The generative module allows imputing the missing modalities in the latent space given the others, when necessary. This task is possible thanks to the injection in the complete framework of four pre-trained generators derived from two different cGANs [42]. The first, the sMRI-SNPs-cGAN allows imputing the latent genetics features transferring the knowledge from the latent sMRI one and vice-versa, and the second, the sMRI-fMRI-cGAN, transfers the knowledge again from the latent sMRI features to generate the respective rs-fMRI ones, and vice-versa. An architecture fMRI-SNPs-cGAN was not developed since the number of subjects that shared both modalities was not sufficient.
The cGANs were built and trained prior to the full framework. In detail, to obtain the pretrained generators, the first step required to obtain the channels’ latent representations (i.e., the $100$ features vector for each modality) needed to subsequently train the cGANs. This was necessary since, as already highlighted, the proposed framework performs the data generation directly in the latent space and not in the original input space. Since the subset of subjects having all three modalities was not numerous enough to train a classification model, two separate bi-modal models were developed, one having sMRI and SNPs (sMRI-SNPs-NN) as input channels and the other having sMRI and rs-fMRI as input channels (sMRI-fMRI-NN). Fig. 2a shows the schematic representation of the sMRI-fMRI-NN, as a reference example. In order to obtain the same latent features as the complete framework, the feature reduction for each channel and the data fusion & classification modules of these bi-modal models were the same as the ones adopted in the complete framework, and described in the previous paragraph. Once the latent vectors were obtained, they were given as input to two cGANs, namely sMRI-fMRI-cGAN and sMRI-SNPs-cGAN, one for each bi-modal model, and hence for each information transfer, from sMRI to rs-fMRI and from sMRI to SNPs, respectively. Fig. 2b shows the sample sMRI-fMRI-cGAN architecture used to generate sMRI from rs-fMRI and vice-versa. The generators are composed of six FCLs (number of nodes: $256$, $512$, $512$, $1024$, $512$, and $100$, respectively) activated by a ReLU activation function. The discriminators consisted of four FCLs (number of nodes: $256$, $128$, $64$, and $1$, respectively) activated by the LeakyRelu function [49], alternated with three dropout layers (dropout probability: $0.3$). The same architecture was considered for the sMRI-SNPs-cGAN.
Figure 2: Schematic representation of a sample bi-modal framework (sMRI-fMRI-NN) followed by the respective cGAN (sMRI-fMRI-cGAN). (a) The sMRI-fMRI-NN is composed of two different modules: feature reduction module, equal to the respective module for each channel in the full framework, and data fusion & classification module where the latent features are fused together and then classified. (b) After training, the latent features of both modalities obtained from the feature reduction module were extracted for each subject and given as input for the training of the associated cGAN, whose loss is described in (c) and (d).
@@
The data fusion & classification module consists of a fusion layer and a classifier. In this module, the latent features obtained either from the data reduction module or the data generation module (for the subjects with missing modalities) are fused together through vector concatenation, resulting in $300$ features. A Multilayer Perceptron (MLP) composed of three FCLs (number of nodes: $150$, $75$, and $2$, respectively) activated by the ReLU function was then used for classification. Softmax was used as activation function for the output layer, allowing to obtain the classification probability associated with the input data.
@@ 2.5 Training scheme
Figure 3: Framework training scheme. The training was performed in two steps: Step 1 involved the training of two bi-modal models (sMRI-fMRI-NN and sMRI-SNPs-NN, Step 1a) and the respective cGANs (sMRI-fMRI-cGAN and sMRI-SNPs-cGAN, Step 1b); Step 2 was needed for the training of the complete multimodal framework.
Figure 3 illustrates the training workflow of the entire framework. As already specified, the training phase was not performed end-to-end but was divided into two steps, necessary for the correct training of the generation module: (i) Step 1: Training of bi-modal models and cGANs; (ii) Step 2: Training of the full multimodal framework.
Hyperparameter tuning was performed empirically. The number of layers, channels, nodes, and the filters’ size were tuned considering the single-modality architectures for the classification (Task 1) before training the multimodal framework. The weights of the CNNs and classifier, used in both bi-modal models and the final pipeline, were initialed as follows: the 3D CNN for sMRI, the 1D CNN for SNPs, and the classifier were initialized using the Xavier Uniform distribution, while the weights of the 1D CNN for rs-fMRI were initialized using the Xavier Normal distribution. The weights of the generators and discriminators were initialized following a Xavier Normal distribution [50] for both the sMRI-fMRI-cGAN and sMRI-SNPs-cGAN. The weights initialization was independent between Step 1 and Step 2.
@@
In the training Step 1 the two sMRI-fMRI-NN and sMRI-SNPs-NN bi-modal models were trained for the AD and CN classification (Step 1a, Fig. 3) in order to subsequently extract the latent features to be fed to the two cGANs, sMRI-fMRI-cGAN and sMRI-SNPs-cGAN(Step 1b, Fig. 3). More in-depth, the bi-model models and subsequently the cGANs were trained relying on the AD and CN subjects sharing respectively the sMRI and rs-fMRI acquisition and the sMRI and SNPs acquisition, presented in Table 1. A $10$-folds stratified Cross Validation (CV) procedure for a total of $40$ epochs, Adam optimizer [51] with a learning rate of $0.0001$, and mini-batch technique considering a batch size of $14$ and $9$ were selected respectively for the training of the sMRI-fMRI-NN and the sMRI-SNPs-NN. Weighted Cross Entropy (CE) was considered as the loss function. Differently from the classical CE, it allows computing a weight guaranteeing higher value to the less numerous class, to address the unbalanced classes issue. The best model in terms of validation accuracy was retained for both the NNs and was used to obtain the $100$ latent features for each subject and for each modality. In detail, the sMRI-fMRI-NN and the sMRI-SNPs-NN achieved an average validation accuracy of $0.890\pm 0.03$ and $0.864\pm 0.03$, and then the best models in terms of accuracy for both bi-modal models were retained for extracting the $100$ latent features vectors obtained at the end of each feature reduction module for each subject.
@@
The newly obtained feature set was then used to train the two cGANs. A mini-batch technique, keeping the same subjects in each batch as the bi-modal models, was used for both the sMRI-fMRI-cGAN and the sMRI-SNPs-cGAN training. Adam optimizer with a learning rate of $0.0001$ was used to train the generators for a total of $600$ epochs. As in [42], the adversarial losses were considered as optimization targets for each cGAN and are represented in Fig. 2c. In detail, considering the sMRI-fMRI-cGAN as reference, given $s$ and $f$ the latent feature vector for the sMRI and rs-fMRI, $G_{fs}$ and $G_{sf}$ the generators to obtain $s$ from $f$ and vice-versa and $D_{s}$ and $D_{f}$ the discriminators to distinguish the $s$ from the generated $\hat{s}$ and the $f$ from the generated $\hat{f}$, respectively, the adversarial loss penalizes the $D_{s}$ and $D_{f}$ errors in discriminating between $\hat{s}=G_{fs}(f)$ and $s$, and $\hat{f}=G_{sf}(s)$ and $f$, respectively. The major issue arising when considering this loss solely is that with a large enough capacity, the generators $G_{fs}$ and $G_{sf}$ can map the same set of input images to any random permutation of images in the target domain (latent features in our case) [42]. The cycle consistency losses were hence introduced to limit the space of possible mapping functions by also penalizing the difference between $f$ and $G_{sf}(G_{fs}(f))$, and between $s$ and $G_{fs}(G_{sf}(s))$, respectively, hence completing the cycle (Fig. 2d).
The four pre-trained generators, two for the sMRI-fMRI-cGAN and two for the sMRI-SNPs-cGAN, were then extracted and injected in the latent space of the proposed multimodal framework for the on-line generation of the missing modalities during the full framework training.
Of note, no data leakage was present between the training Step 1 (bi-modal models and cGANs) and the training Step 2 (full multimodal framework) since Step 2 involved end-to-end training of subjects with no missing modalities (i.e., using input data for these subjects only) and generative models trained in Step 1b for deriving latent features for subjects with missing modalities only. Of note, the full cohort of subjects was not used to train the bi-modal models and corresponding cGANs (Step 1a and 1b), only relying on those subjects sharing the modalities under analysis (sMRI and rs-fMRI for sMRI-fMRI-NN/cGAN, and sMRI and SNPs for sMRI-SNPs-NN/cGAN).
The cycle consistency loss was used in both sMRI-fMRI-cGAN and sMRI-SNPs-cGAN in order to assess the performance of the generators in imputing missing modalities using the Mean Absolute Error (MAE).
@@
The full multimodal framework including the three input channels, sMRI, rs-fMRI, and SNPs, was trained for classification Task 1, AD detection, relying on the complete AD and CN data cohort.
Subjects were split into training and testing sets, respectively, including the $80\%$-$20\%$ of the study cohort. The testing set was kept unseen until the last testing phase. A stratified $10$-fold CV procedure was applied to the training set. A mini-batch strategy ($12$ subjects in each batch) was adopted during training. Adam optimizer with a learning rate of $0.0001$ was chosen and the model was trained for a total of $70$ epochs. Weighted CE was considered as the loss function. During this training phase, the four pre-trained cGAN generators’ weights were kept frozen. Indeed, during the backpropagation phase, only the weights of the feature reduction module and the data fusion and classification module were updated.

### v3

@@ 2.2 Preprocessing and feature engineering
The sMRI volumes preprocessing included tissue segmentation in Gray Matter (GM), White matter, and Cerebrospinal fluid (CSF) using the modulated normalization algorithm. Only the GM volume was considered for this study, to which a smoothing using a Gaussian kernel (FWHM = $6$mm) was applied. The full preprocessed GM volume was used as input for the sMRI channel resulting in an input size of $121\times 145\times 121$ for each subject. The rs-fMRI data was preprocessed using the statistical parametric mapping toolbox (SPM12, http://www.fil.ion.ucl.ac.uk/spm/) including rigid body motion correction to correct subject head motion, slice-timing correction, warping to the standard MNI space using the EPI template, resampling to ${(3mm)}^{3}$ isotropic voxels, and smoothing using a Gaussian kernel (FWHM = 6mm), following the preprocessing proposed in [58]. Fifty-three independent components (ICs) covering the whole brain were extracted using spatially constrained ICA with the Neuromark_fMRI_1.01.01.01.0 template (available in the GIFT software; http://trendscenter.org/software/gift). For each subject, a correlation matrix was then created computing the Pearson correlation between the ICs time courses, resulting in a $53$x53535353 static functional network connectivity (sFNC) matrix. This was divided into $7$ RSNs, named: (i) Sub-cortical network (SC), (ii) Auditory network, (iii) Sensorimotor network (SM), (iv) Visual network (VI), (v) Cognitive-control network (CC), (vi) Default-mode network (DM), and (vii) Cerebellar network (CB) (please, refer to Supplementary Table 2 for more information). The upper triangular matrix was then vectorized, resulting in an input vector of 1378 features for each subject. Quality control of MRI images is detailed in the supplementary materials section "Materials and Methods", subsection "Preprocessing quality control of MRI".
@@
Moving to the genomics data, pre-imputation QC was performed to remove Single Nucleotide Polymorphisms (SNPs) with minor allele frequency (MAF) $<0.05$, call rate $<0.98$, and Hardy Weinberg Equilibrium $<10^{-3}$ (details in https://www.synapse.org/#!Synapse:syn2290704/wiki/64710). Imputation was performed with the same reference panel. Only the SNPs with imputation $r^{2}>0.4$ were included. Linkage disequilibrium (LD)-based pruning with $r^{2}=0.8$ in a window of $50$kb was applied, yielding $445838$ SNPs for further analyses [59]. A feature selection leveraging on the genome-wide association study (GWAS) on AD conducted by [17] was carried on to include all the relevant SNPs having GWAS p-values less than $1$e-04. A total of $565$ SNPs was selected and a value between 0 and 2 was assigned based on the presence of mutated alleles. In detail, defining wild and mutated alleles respectively as W and M, a score of 00 means wild homozygous alleles (W/W), a score of $1$ indicates heterozygous alleles (W/M), and a score of $2$ defines mutated homozygous alleles (M/M).
@@
Figure 1: The proposed framework that integrates sMRI, fMRI, and SNPs is articulated in three modules: (i) A feature reduction module where the latent representations are extracted for the input data; (ii) A generative module where, if necessary, the missing modalities are imputed; and (iii) A data fusion & classification module where the latent features are fused and then classified. Finally, post-hoc XAI analysis is performed relying on the IG method for feature attribution derivation.
2.3 Framework architecture
The proposed framework builds on our previous work [35] and employed for addressing two classification tasks: Task 1, CN-AD, and Task 2, MCInc-MCIc, is shown in Figure 1. In detail, the architecture for disease detection consists of three modules: (i) Feature reduction module, which performs a CNN-based feature extraction to derive a lower dimensionality latent space, separately for each input channel; (ii) Generative module that, in the eventuality of missing modalities, actuates a generative process in the latent space transferring the knowledge from one domain to another; and (iii) Data fusion & classification module that fuses the latent features obtained for the three modalities and then performs the classification. Post-hoc interpretability analysis was then carried on in order to retrieve features’ contributions to the classification task. The three modules will be detailed in the following paragraphs.
@@ 2.3 Framework architecture
The feature reduction module consists of three different CNNs, one for each input channel, leading to a latent low dimensionality representation consisting of $100$ latent features for each channel. The sMRI channel was analyzed through a 3D CNN defined by six convolutional layers (filter sizes of $3\times 3\times 3$ for the first four convolutional layers and $2\times 2\times 2$ for the last two, number of filters: $64$, $64$, $64$, $128$, $128$, and $128$) and three max-pooling layer, followed by four fully-connected layers (FCLs) (number of nodes: $1536$, $768$, $384$, and $192$). The rs-fMRI channel was analyzed through a $1$D CNN, consisting of four convolutional layers (filter sizes $5$, number of filters: $64$, $128$, $128$, and $128$), two max-pooling layers, and two FCLs (number of nodes: $384$ and $100$). Finally, for genetic data, a $1$D CNN was employed, enclosing three convolutional blocks, each including a sequence of convolutional layers (filter sizes $3$, number of filters: $64$, $64$, and $128$), batch-normalization and max-pooling, followed by three FCLs (number of nodes: $1024$, $512$, and $128$). The ReLU activation function was used for all the layers of these three CNNs.
@@
The generative module allows imputing the missing modalities in the latent space given the others, when necessary. This task is possible thanks to the injection in the complete framework of four pre-trained generators derived from two different cGANs [53]. The first, the sMRI-SNPs-cGAN allows imputing the latent genetics features transferring the knowledge from the latent sMRI one and vice-versa, and the second, the sMRI-fMRI-cGAN, transfers the knowledge again from the latent sMRI features to generate the respective rs-fMRI ones, and vice-versa. The scarcity of subjects sharing all modalities impeded the development of an fMRI-SNPs-cGAN model.
The cGANs were built and trained prior to the full framework. In detail, to obtain the pretrained generators, the first step required to obtain the channels’ latent representations (i.e., the $100$ features vector for each modality) needed to subsequently train the cGANs. This was necessary since, as already highlighted, the proposed framework performs the data generation directly in the latent space and not in the original input space. Since the subset of subjects having all three modalities was not numerous enough to train a classification model, two separate bi-modal models were developed, one having sMRI and SNPs (sMRI-SNPs-NN) as input channels and the other having sMRI and rs-fMRI as input channels (sMRI-fMRI-NN). Figure 2a shows the schematic representation of the sMRI-fMRI-NN, as a reference example. The feature reduction for each channel and the data fusion & classification modules of these bi-modal models were the same as the ones adopted in the complete framework and described in the previous paragraph. Once the latent vectors were obtained, they were given as input to two cGANs, namely sMRI-fMRI-cGAN and sMRI-SNPs-cGAN, one for each bi-modal model, and hence for each information transfer, from sMRI to rs-fMRI and from sMRI to SNPs, respectively. Figure 2b shows the sample sMRI-fMRI-cGAN architecture used to generate sMRI from rs-fMRI and vice-versa. The generators are composed of six FCLs (number of nodes: $256$, $512$, $512$, $1024$, $512$, and $100$) activated by a ReLU activation function. The discriminators consisted of four FCLs (number of nodes: $256$, $128$, $64$, and $1$) activated by the LeakyRelu function, alternated with three dropout layers (dropout probability: $0.3$). The same architecture was considered for the sMRI-SNPs-cGAN.
Figure 2: The bi-modal model sMRI-fMRI-NN/cGAN. a) The sMRI-fMRI-NN is composed of feature reduction and data fusion & classification modules (equal architectures of the full model). b) After training, the feature reduction module extracts the latent features of both MRI used for training the sMRI-fMRI-cGAN, whose losses are described in c) and d).
@@
The data fusion & classification module consists of a fusion layer and a classifier. The latent features obtained either from the data reduction module or the data generation module (for the subjects with missing modalities) are fused through vector concatenation, resulting in $300$ features. A multilayer perceptron composed of three FCLs (number of nodes: $150$, $75$, and $2$) activated by the ReLU function was then used for classification. Softmax activated the output layer.
@@ 2.4 Training scheme
Figure 3: The framework training is performed in two steps: Step 1 involves the training of sMRI-fMRI-NN and sMRI-SNPs-NN (Step 1a) and the respective sMRI-fMRI-cGAN and sMRI-SNPs-cGAN (Step 1b); while in Step 2 the multimodal framework is trained.
Figure 3 illustrates the training workflow of the entire framework. As already specified, the training phase was not performed end-to-end but was divided into two steps, each one with end-to-end training, necessary for the correct training of the generation module: (i) Step 1: Training of bi-modal models and cGANs, and (ii) Step 2: Training of the full framework.
Hyperparameter tuning was performed through a grid search. The number of layers, channels, nodes, and the filters’ size were tuned considering the single-modality architectures for the classification (Task 1) before training the multimodal framework.
@@
In the training Step 1 the two sMRI-fMRI-NN and sMRI-SNPs-NN bi-modal models were trained for the AD versus CN classification (Step 1a, Figure 3) in order to subsequently extract the latent features to be fed to the two cGANs, sMRI-fMRI-cGAN and sMRI-SNPs-cGAN (Step 1b, Figure 3). More in-depth, the bi-model models and subsequently the cGANs were trained relying on the AD and CN subjects sharing respectively the sMRI and rs-fMRI acquisition and the sMRI and SNPs acquisition, presented in Table 1. A $10$-folds stratified Cross Validation (CV) procedure for a total of $40$ epochs, Adam optimizer [60] with a learning rate of $0.0001$, and mini-batch technique considering a batch size of $14$ and $9$ were selected, respectively, for the training of the sMRI-fMRI-NN and the sMRI-SNPs-NN. Weighted Cross Entropy (CE) was used as the loss function. The sMRI-fMRI-NN and the sMRI-SNPs-NN achieved an average validation accuracy of $0.89\pm 0.03$ and $0.864\pm 0.03$, respectively. The best model in terms of validation accuracy was retained for both the NNs and used to extract the $100$ latent features for each subject and for each modality. The newly obtained feature set was used to train the two cGANs. A mini-batch technique, keeping the same subjects in each batch as the bi-modal models, was used for both the sMRI-fMRI-cGAN and the sMRI-SNPs-cGAN training. Adam optimizer with a learning rate of $0.0001$ was used to train the generators for a total of $600$ epochs. As in [53], the adversarial losses were considered as optimization targets for each cGAN as in Figure 2c. In detail, considering the sMRI-fMRI-cGAN as reference, given $s$ and $f$ the latent feature vectors for the sMRI and rs-fMRI, $G_{fs}$ and $G_{sf}$ the generators to obtain $s$ from $f$ and vice-versa and $D_{s}$ and $D_{f}$ the discriminators to distinguish the $s$ from the generated $\hat{s}$ and the $f$ from the generated $\hat{f}$, respectively, the adversarial loss penalizes the $D_{s}$ and $D_{f}$ errors in discriminating between $\hat{s}=G_{fs}(f)$ and $s$, and $\hat{f}=G_{sf}(s)$ and $f$, respectively. The major issue arising when considering this loss solely is that with a large enough capacity, the generators $G_{fs}$ and $G_{sf}$ can map the same set of input images to any random permutation of images in the target domain (latent features in our case) [53]. The cycle consistency losses were hence introduced to limit the space of possible mapping functions by also penalizing the difference between $f$ and $G_{sf}(G_{fs}(f))$, and between $s$ and $G_{fs}(G_{sf}(s))$, respectively, hence completing the cycle (Figure 2d). The cycle consistency loss was used in both sMRI-fMRI-cGAN and sMRI-SNPs-cGAN in order to assess the performance of the generators in imputing missing modalities using the Mean Absolute Error (MAE). The four pre-trained generators, two for the sMRI-fMRI-cGAN and two for the sMRI-SNPs-cGAN, were then extracted and injected in the latent space of the proposed multimodal framework for the online generation of the missing modalities during the full framework training.
@@
The full multimodal framework including the three input channels, sMRI, rs-fMRI, and SNPs, was trained for classification Task 1, AD detection, relying on the complete AD and CN data cohort. Subjects were split into training and testing sets, respectively, including the $80\%$-$20\%$ of the study cohort. The testing set was kept unseen until the last testing phase. A stratified $10$-fold CV procedure was applied to the training set. A mini-batch strategy ($12$ subjects per batch) was adopted during training. Adam optimizer with a learning rate of $0.0001$ was chosen and the model was trained for a total of $70$ epochs. Weighted CE was considered as the loss function. During this training phase, the four pre-trained cGAN generators’ weights were kept frozen. Indeed, during backpropagation, only the weights of the feature reduction module and the data fusion and classification module were updated. Additional information about the training procedure is described in the supplementary materials "Materials and Methods" and subsection "Training scheme additional information".